HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025, Page 4 of 7  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  75   Wed Apr 2 10:31:15 2025 Michael, AlexAcceleratorThe beam is back from 10:10     
  74   Wed Apr 2 10:05:06 2025 Michael, Alex, YuriAcceleratorNo beam since 09:59The SIS kicker has failed again.   
  73   Wed Apr 2 10:02:40 2025 Sergey + YuriAcceleratorESR TimeoutAt around 9:40, ESR timeout changed from 10
s to 30 s
Triggering only on real injections.
  
  72   Wed Apr 2 09:28:59 2025 Michael, AlexGeneralEmpty injections in the ESRSince the FRS started to take the beam again,
the ESR gets 1/10 shots from SIS. But, the
ESR seems to request about twice more frequently
  
  71   Wed Apr 2 08:47:27 2025 Michael, AlexGeneralThe beam is backThe beam is back. But, the experts have found
a problem with one of the modules of the
SIS kicker magnet. And, currently they have
  
  70   Wed Apr 2 08:04:47 2025 Michael, AlexGeneralNo beam The SIS kicker has failed again at 08:01.
Since this has happened already few times,
the experts from the on-call duty service
  
  69   Wed Apr 2 07:41:10 2025 Michael, AlexGeneralFrequency shift in SchottkyAfter the beam  was restored at 07:25, we
think that the frequency peak in the 244
MHz RSA2 has jumped again (closer to the
  
  68   Wed Apr 2 07:39:41 2025 Michael, AlexGeneralThe beam is back at 07:35    
  67   Wed Apr 2 07:33:59 2025 Michael, AlexGeneralBeam gone again7:29, beam gone again, operators say they
need another 5 minutes.
  
  66   Wed Apr 2 07:25:15 2025 Carlo, Chris, Alex, Michael GeneralBeam back07:20  we have beam back.   
  65   Wed Apr 2 07:16:33 2025 Carlo, Chris, Alex, MichaelDAQESR not requesting7:05 ESR not requesting new pulses from the
SIS. 
Operators are trying to fix the problem.
  
  64   Wed Apr 2 05:48:31 2025 Chris, Carlo, ZacGeneralBeam returnedBeam was returned to the ESR at 05:11.

HKR said there had been some problem with
  
  63   Wed Apr 2 04:59:50 2025 Zac, Chris, CarloGeneralReceiving injections but not seeing anything on RSAsStarting at around 04:09 we don't see anything
on RSAs. RSAs are still triggering and writing
files, but only see noise. Lassie indicates
 2025-04-02_05-18-32-212_tcl1146.png2025-04-02_05-21-50-973_tcl1146.png 
  62   Wed Apr 2 04:33:57 2025 Carlo, Zac, ChrisDAQScreenshots Lassiescreenshots of Lassie  2025-04-02_04-28-59-944_tcl1064.png2025-04-02_03-38-54-877_tcl1064.png 
  61   Wed Apr 2 02:51:31 2025 Carlo, Zac, ChrisDAQBeam is back02:44 beam is back   
  60   Wed Apr 2 01:46:52 2025 Chris, Carlo, ZacGeneralBeam lossHKR just called to say a power supply to the
SIS18 extraction kicker has failed.
The service person has been called and is
  
  59   Wed Apr 2 00:59:43 2025 Zac, Carlo, ChrisAcceleratorLost the beamBeam was lost from 00:44 until 00:59   
  58   Wed Apr 2 00:08:57 2025 Ruijiu Chen,ShahabAnalysisthe peak corresponding to 194Pb⁸²⁺ was mixed with signals from some unknown ions.I examined the spectrum of 194Pb at the 211
harmonic. It was observed that the peak corresponding
to 194Pb82+ was mixed with signals 
 photo_2025-04-02_00-12-38.jpgphoto_2025-04-02_00-12-43.jpg 
  57   Tue Apr 1 23:56:18 2025 Shahab, Carlo, Zac, Yue, MichaelDAQChange of RSA01 settings for new harmonic=212.23:55 stop the RSAs. 

410:
  
  56   Tue Apr 1 23:23:18 2025 C GriffinGeneralPLEIADES removedLeakage currents were steadily increasing
in the PLEIADES detectors so we decided to
stop taking data, move both the detectors
  
ELOG V3.1.5-fc6679b